OpenStudy (cutebunny12):

i will medal who ever gives me the answer When solving negative one over three.(x − 15) = −4, what is the correct sequence of operations? Multiply each side by negative one over three , subtract 15 from each side Multiply each side by −3, add 15 to each side Multiply each side by negative one over three , add 15 to each side Multiply each side by 3, subtract 15 from each side

OpenStudy (benlindquist):

easy, what do you say

OpenStudy (cutebunny12):

i dont know

OpenStudy (benlindquist):

guess

OpenStudy (cutebunny12):

b

OpenStudy (cutebunny12):

is it B

OpenStudy (benlindquist):

no

OpenStudy (cutebunny12):

um c

OpenStudy (cutebunny12):

wait no its D

OpenStudy (anonymous):

-1/3(x-15)=-4 x3 x3 -x+15= -12 -15 -15 -x = -27 x=27 D is correct

OpenStudy (cutebunny12):

thank you

OpenStudy (anonymous):

@cutebunny12 no problem :)